In a certain government office, there are 400 employee, there are 150 men, 276 University graduates, 212 married persons, 94 male University graduates, 151 married university graduates, 119 married men, 72 married male graduates. Find the number of single women who are not University graduates.
1
Expert's answer
2020-04-13T10:12:28-0400

Among 400 employee there are 150 men, so there are 400 - 150 = 250 women.

Among 276 University graduates there are 94 male University graduates, so there are 276 - 94 = 182 female University graduates, then there are 250 - 182 = 68 female non graduate.

Among 212 married persons there are 119 married men, so there are 212 - 119 = 93 married women, then there are 250 - 93 = 157 single women.

Among 151 married university graduates, there are 72 married male graduates, so there are 151 - 72 = 79 married female graduates.

The number of single female non graduates is: 250 - (182 + 93 - 79) = 54 women.